# 4.2cm4.2cm3.6cm3cm

### v1

@@ 3. $2$-local derivations on $\mathcal{L}(\Gamma)$
$\Delta(x)$ $=$ $\Delta_{L_{0,i},x}(x)$
$=$ $[a_{0,j}(L_{0,i},x)L_{0,j}+b_{0,j}(L_{0,i},x)H_{0,j},x]+\lambda_{3}(L_{0,i},x)D_{b}(x)$
$=$ $\sum_{\begin{subarray}{c}\gamma\in\Gamma k\in\mathbb{Z}\end{subarray}}(-\gamma m_{\gamma,k}a_{0,j}L_{\gamma,j+k}-\gamma n_{\gamma,k}a_{0,j}(L_{0,i},x)H_{\gamma,j+k})+n_{\gamma,k}b\lambda_{3}(L_{0,i},x)(H_{\gamma,k})$,
$\Delta(x)$ $=$ $\Delta_{L_{\alpha,i},x}(x)$
$=$ $[a_{\alpha,j}(L_{\alpha,i},x)L_{\alpha,j}+b_{0,j}(L_{\alpha,i},x)H_{0,j},x]+(1-\delta_{\alpha,0})\lambda_{1}(L_{\alpha,i},x)D_{f}(x)$
$+(1-\delta_{\alpha,0})\lambda_{2}(L_{\alpha,i},x)D_{g}(x)+\lambda_{3}(L_{\alpha,i},x)D_{b}(x)+(1-\delta_{i,0})\lambda_{4}(L_{\alpha,i},x)D^{\rho}(x)$
$=$ $\sum_{\begin{subarray}{c}\gamma\in\Gamma\\ k\in\mathbb{Z}\end{subarray}}((\alpha-\gamma)m_{\gamma,k}a_{\alpha,j}(L_{\alpha,i},x)L_{\alpha+\gamma,j+k}-\gamma n_{\gamma,k}a_{\alpha,j}(L_{\alpha,i},x)H_{\alpha+\gamma,j+k})$
$+f_{\gamma}(1-\delta_{\alpha,0})\lambda_{1}(L_{\alpha,i},x)(m_{\gamma,k}L_{\gamma,k}+n_{\gamma,k}H_{\gamma,k})+g_{\alpha}(1-\delta_{\alpha,0})\lambda_{2}(L_{\alpha,i},x)(m_{\gamma,k}L_{\gamma,k})$
$+b\lambda_{3}(L_{\alpha,i},x)(n_{\gamma,k}H_{\gamma,k})+k(1-\delta_{i,0})\lambda_{4}(L_{\alpha,i},x)(m_{\gamma,k}L_{\gamma,k-1}+n_{\gamma,k}H_{\gamma,k-1})$.
If there exists $m_{\alpha,i}\neq 0$, taking sufficiently many $\alpha\in\Gamma,\,i\in\mathbb{Z}$, we have $a_{\alpha,j}(L_{\alpha,i},x)=a_{0,j}(L_{\alpha,i},x)=0,\,i,j\in\mathbb{Z}$. Similarly $\lambda_{3}(L_{0,i},x)=\lambda_{3}(L_{\alpha,i},x)=0$, and $\lambda_{1}(L_{\alpha,i},x)=\lambda_{2}(L_{\alpha,i},x)=0(\alpha\neq 0),\,\lambda_{4}(L_{\alpha,i},x)=0(i\neq 0)$. When $\alpha=0$, we have $f_{\alpha}(1-\delta_{\alpha,0})$ and $g_{\alpha}(1-\delta_{\alpha,0})=0$. When $i=0$, we have $i(1-\delta_{i,0})=0$. ∎
@@
$\Delta(L_{0,i})$ $=$ $\Delta_{L_{0,i},L_{1,j}}(L_{0,i})$,
$\Delta(L_{1,j})$ $=$ $\Delta_{L_{0,i},L_{1,j}}(L_{1,j})$.
@@ 4. Biderivations on $\mathcal{L}(\Gamma)$
Recall the definitions of biderivations and inner biderivations. Let $L$ be an algebra and $f$ be a bilinear map on $L\times L\rightarrow L$. If it is a derivation with respect to both components, meaning that for all $x\in L$, there exists a linear map from $L$ to itself $\phi_{x},\psi_{x}$, and $\phi_{x}=f(x,\cdot),\psi_{x}=f(\cdot,x)$ are derivations on $L$, satisfying that
@@
Proof.
Since [,][,][ , ] is bilinear, we only need to verify
$L_{\alpha,i}$ $=$ $\frac{1}{\alpha}[L_{\alpha-\beta,i-j},L_{\beta,j}]\quad(\alpha\neq 0,\alpha,\beta\in\Gamma,i,j\in\mathbb{Z}),$
$L_{0,i}$ $=$ $\frac{1}{2}[L_{1,i-j},L_{-1,j}]\quad(\alpha=0,i,j\in\mathbb{Z}),$
$H_{\alpha,i}$ $=$ $-\frac{1}{\beta}[L_{\alpha-\beta,i-j},H_{\beta,j}]\quad(\beta\neq 0,\alpha,\beta\in\Gamma,i,j\in\mathbb{Z}),$
$H_{0,i}$ $=$ $[L_{1,i-j},H_{-1,j}]\quad(\beta=0,i,j\in\mathbb{Z}).$
Then (Γ)=[$\mathcal{L}(\Gamma)=[\mathcal{L}(\Gamma),\mathcal{L}(\Gamma)]$. ∎
@@
Let $f$ be biderivation on $\mathcal{L}(\Gamma)$.
(i) There exists $\lambda\in\mathbb{F}$ such that $f(L_{0,i},L_{m,j})\equiv\lambda[L_{0,i},L_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m\in\Gamma,i,j\in\mathbb{Z}.$
If $m=0$, then $[L_{0,i},L_{0,j}]=0$, by Lemma 15 we conclude that
@@
$[[L_{0,i},L_{m,j}],f(L_{0,i},L_{m,j})]=0.$
Since
$[-mL_{m,i+j},a_{n}L_{n,i+j}+b_{n}H_{n,i+j}]=0,\,n\in\Gamma,i,j\in\mathbb{Z},$
then it follows that
$-ma_{n}(m-n)=0,\quad mnb_{n}=0.$
Furthermore, $a_{n}=0(m\neq n),\,b_{n}=0(n\neq 0)$, we conclude that
$f(L_{0,i},L_{m,j})$$\equiv$ $0(\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m,n\neq 0,$
$f(L_{0,i},L_{m,j})$$\equiv$ $a_{m}L_{m,i+j}(\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m\neq n,n=0.$
Thus $f(L_{0,i},L_{m,j})\equiv\lambda[L_{0,i},L_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m\in\Gamma,i,j\in\mathbb{Z}$.
(ii)$f(L_{0,i},H_{m,j})\equiv\lambda[L_{0,i},H_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m\in\Gamma,i,j\in\mathbb{Z}$.
If $m=0$, $H_{0,j}\in C(\mathcal{L}(\Gamma))$, then $[L_{0,i},H_{0,j}]=0$, it follows that $f(L_{0,i},H_{0,j})\in C(\mathcal{L}(\Gamma))$.
@@
$[[L_{0,j},H_{m,j}],f(L_{0,j},H_{m,j})]=0,$
then we have
$[-mH_{m,i+j},a_{n}L_{n,i+j}+b_{n}H_{n,i+j}]=0,\,n\in\Gamma,i,j\in\mathbb{Z}.$
We conclude $-m^{2}a_{n}=0$, thus $a_{n}=0$, then
@@
$-nb_{n}H_{n+1,2(i+j)}=-m\lambda H_{m+1,2(i+j)}.$
We have $m=n$, which gives $f(L_{0,i},H_{m,j})\equiv\lambda[L_{0,i},H_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m\in\Gamma,i,j\in\mathbb{Z}$.
(iii)$f(L_{0,i},x)\equiv\lambda[L_{0,i},x](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,x\in\mathcal{L}(\Gamma),i,j\in\mathbb{Z}$.
@@
We have
$f(L_{0,i},x)$ $=$ $\sum_{\begin{subarray}{c}m\in\Gamma\\ i,j\in\mathbb{Z}\end{subarray}}\left\{a_{m}f(L_{0,i},L_{m,j})+b_{m}f(L_{0,i},H_{m,j})\right\}$
$=$ $\sum_{\begin{subarray}{c}m\in\Gamma\\ i,j\in\mathbb{Z}\end{subarray}}\left\{a_{m}\lambda[L_{0,i},L_{m,j}]+b_{m}\lambda[L_{0,i},H_{m,j}]\right\}(\mathrm{mod}\mathbb{F}H_{0,i+j})$
@@
which gives $f(L_{0,i},x)\equiv\lambda[L_{0,i},x](\mathrm{mod}\mathbb{F}H_{0,i+j}),x\in\mathcal{L}(\Gamma),i,j\in\mathbb{Z}$.
(iv)$f(x,y)\equiv\lambda[x,y](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,x,y\in\mathcal{L}(\Gamma),i,j\in\mathbb{Z}$.
@@
$[f(x,y),[L_{0,i},z]]=[[x,y],f(L_{0,i},z)],\,z\in\mathcal{L}(\Gamma),$
we have
$[f(x,y)-\lambda[x,y],[L_{0,i},z]]=0,\,z\in\mathcal{L}(\Gamma),$
which gives $f(x,y)-\lambda[x,y]\in C(\mathcal{L}(\Gamma))$, thus $f(x,y)\equiv\lambda[x,y](\mathrm{mod}\mathbb{F}H_{0,i+j})$.
(v)$f(x,y)\equiv\lambda[x,y],\,x,y\in\mathcal{L}(\Gamma)$.
@@
where $\alpha$ is a bilinear map on $\mathcal{L}(\Gamma)\times\mathcal{L}(\Gamma)\rightarrow\mathbb{F}$. By Lemma 13
@@
We have
$\alpha([x,y],z)=0,\,x,y,z\in\mathcal{L}(\Gamma).$
Since $\mathcal{L}(\Gamma)$ is perfect, we conclude $\alpha(x,y)=0,\,x,y\in\mathcal{L}(\Gamma)$. Finally, we have that $f(x,y)\equiv\lambda[x,y],\,\forall x,y\in\mathcal{L}(\Gamma)$. That is, $f$ is an inner derivation on $\mathcal{L}(\Gamma)$. ∎
@@ 4.2. Applications
Recall the definition of linear commuting maps. If a linear map $\phi$ on Lie algebra $L$, which subjects to $[\phi(x),x]=0,\,\forall x\in L$, then $\phi$ called a linear commuting map on $L$. If $f(x,y)=[\phi(x),y]=[x,\phi(y)]$, then $f$ is a biderivation on $L$. (see [12])
@@
Any linear map on $\mathcal{L}(\Gamma)$ is commuting if and only if there exists $\lambda\in\mathbb{F}$ and a linear map $\tau:\,\mathcal{L}(\Gamma)\rightarrow C(\mathcal{L}(\Gamma))$ such that $\phi(x)=\lambda x+\tau(x),\,x\in\mathcal{L}(\Gamma)$.
@@
Conversely, let $f(x,y)=[\phi(x),y],\,x,y\in\mathcal{L}(\Gamma)$. Then $f$ is a biderivation on $\mathcal{L}(\Gamma)$. By Theorem 15, we have
$f(x,y)=[\phi(x),y]=\lambda[x,y].$
Since $\phi(x)-\lambda x\in C(\mathcal{L}(\Gamma))$, then we have a linear map $\tau:\mathcal{L}(\Gamma)\rightarrow C(\mathcal{L}(\Gamma))$ such that $\phi(x)=\lambda x+\tau(x),\,x\in\mathcal{L}(\Gamma)$. ∎
[…]
@@
$x\cdot[y,z]=[x\cdot y,z]+[y,x\cdot z],\,x,y,z\in L.$
We also call $(\mathcal{L}(\Gamma),[,],\cdot)$ is a commutative $post$-Lie algebra (see [13]).
@@
(see, [13]) Let $(L,[,],\cdot)$ be a commutative $post$-Lie algebra. If $f:L\times L\rightarrow L$ subjects to $f(x,y)=x\cdot y,\,x,y\in L$, then $f$ is a biderivation on $L$.
@@
Any commutative $post$-Lie algebra structure on $\mathcal{L}(\Gamma)$ is trivial.
@@
Assume $(\mathcal{L}(\Gamma),[,],\cdot)$ is a commutative $post$-Lie algebra. By Theorem 15 and Lemma 17, we know there exists $\lambda\in\mathbb{C}$ such that
@@
Thus $02\lambda[x,y]=$ 0, we conclude $\lambda=0$. That is $f(x,y)=x\cdot y=0$. ∎

### v2

@@ 3. $2$-local derivations on $\mathcal{L}(\Gamma)$
$\Delta(x)$ == $\Delta_{L_{0,i},x}(x)$
== $a_{0,j}(L_{0,i},x)L_{0,j}+b_{0,j}(L_{0,i},x)H_{0,j},x]+\lambda_{3}(L_{0,i},x)D_{b}(x)+(1-\delta_{i,0})\lambda_{4}(L_{0,i},x)D^{\rho}(x)$
== $\sum_{\begin{subarray}{c}\gamma\in\Gamma k\in\mathbb{Z}\end{subarray}}(-\gamma m_{\gamma,k}a_{0,j}(L_{0,i},x)L_{\gamma,j+k}-\gamma n_{\gamma,k}a_{0,j}(L_{0,i},x)H_{\gamma,j+k})$
$n_{\gamma,k}b\lambda_{3}(L_{0,i},x)(H_{\gamma,k})+k(1-\delta_{i,0})\lambda_{4}(L_{0,i},x)(m_{\gamma,k}L_{\gamma,k-1}+n_{\gamma,k}H_{\gamma,k-1})$,
$\Delta(x)$ == $\Delta_{L_{\alpha,i},x}(x)$
== $a_{\alpha,j}(L_{\alpha,i},x)L_{\alpha,j}+b_{0,j}(L_{\alpha,i},x)H_{0,j},x]+(1-\delta_{\alpha,0})\lambda_{1}(L_{\alpha,i},x)D_{\phi}(x)$
$(1-\delta_{\alpha,0})\lambda_{2}(L_{\alpha,i},x)D_{g}(x)+\lambda_{3}(L_{\alpha,i},x)D_{b}(x)+(1-\delta_{i,0})\lambda_{4}(L_{\alpha,i},x)D^{\rho}(x)$
== $\sum_{\begin{subarray}{c}\gamma\in\Gamma\\ k\in\mathbb{Z}\end{subarray}}((\alpha-\gamma)m_{\gamma,k}a_{\alpha,j}(L_{\alpha,i},x)L_{\alpha+\gamma,j+k}-\gamma n_{\gamma,k}a_{\alpha,j}(L_{\alpha,i},x)H_{\alpha+\gamma,j+k})$
$\phi(\gamma)(1-\delta_{\alpha,0})\lambda_{1}(L_{\alpha,i},x)(m_{\gamma,k}L_{\gamma,k}+n_{\gamma,k}H_{\gamma,k})+g_{\alpha}(1-\delta_{\alpha,0})\lambda_{2}(L_{\alpha,i},x)(m_{\gamma,k}L_{\gamma,k})$
$b\lambda_{3}(L_{\alpha,i},x)(n_{\gamma,k}H_{\gamma,k})+k(1-\delta_{i,0})\lambda_{4}(L_{\alpha,i},x)(m_{\gamma,k}L_{\gamma,k-1}+n_{\gamma,k}H_{\gamma,k-1})$.
If there exists $m_{\alpha,i}\neq 0$, taking sufficiently many $\alpha\in\Gamma,\,i\in\mathbb{Z}$, then we have $a_{\alpha,j}(L_{\alpha,i},x)=a_{0,j}(L_{\alpha,i},x)=0$ for all $i,j\in\mathbb{Z}$. Similarly $\lambda_{3}(L_{0,i},x)=\lambda_{3}(L_{\alpha,i},x)=0$, and $\lambda_{1}(L_{\alpha,i},x)=\lambda_{2}(L_{\alpha,i},x)=0\,(\alpha\neq 0),\,\lambda_{4}(L_{\alpha,i},x)=0\,(i\neq 0)$. When $\alpha=0$, we have $\phi(\alpha)(1-\delta_{\alpha,0})$ and $g_{\alpha}(1-\delta_{\alpha,0})=0$. When $i=0$, we have $i(1-\delta_{i,0})=0$. ∎
@@
$\Delta(L_{0,i})$ == $\Delta_{L_{0,i},L_{1,j}}(L_{0,i})$,
$\Delta(L_{1,j})$ == $\Delta_{L_{0,i},L_{1,j}}(L_{1,j})$.
@@ 4.1. Biderivations on $\mathcal{L}(\Gamma)$
Recall the definitions of biderivations and inner biderivations. Let $L$ be an algebra and $f$ be a bilinear map on $L\times L\rightarrow L$. If it is a derivation with respect to both components, meaning that for all $x\in L$, there exists linear maps $\phi_{x},\psi_{x}$ from $L$ to itself, and $\phi_{x}=f(x,\cdot),\psi_{x}=f(\cdot,x)$ are derivations on $L$, satisfying that
@@
Let $f$ be biderivation on $\mathcal{L}(\Gamma)$. We proceed by several steps.
(i) First, we show that there exists $\lambda\in\mathbb{F}$ such that $f(L_{0,i},L_{m,j})\equiv\lambda[L_{0,i},L_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j})$ for all $m\in\Gamma$ and $i,j\in\mathbb{Z}.$
If $m=0$, then $L_{0,i},L_{0,j}]=0$, by Lemma 14 we conclude that
@@
$L_{0,i},L_{m,j}],f(L_{0,i},L_{m,j})]=0.$
Since $mL_{m,i+j},\sum_{\begin{subarray}{c}n\in\Gamma\\ i,j\in\mathbb{Z}\end{subarray}}(a_{n}L_{n,i+j}+b_{n}H_{n,i+j})]=0$ for all $n\in\Gamma$ and all $i,j\in\mathbb{Z},$ then it follows that
$\sum_{\begin{subarray}{c}n\in\Gamma\end{subarray}}(-ma_{n}(m-n))=0,\quad\sum_{\begin{subarray}{c}n\in\Gamma\end{subarray}}mnb_{n}=0.$
Furthermore, $a_{n}=0$ for $m\neq n$, $b_{n}=0$ for $n\neq 0$, we conclude that
$f(L_{0,i},L_{m,j})$ $\equiv$ $0(\mathrm{mod}\mathbb{F}H_{0,i+j}),\,n=0,$
$f(L_{0,i},L_{m,j})$ $\equiv$ $a_{m}L_{m,i+j},\,m=n,$
$f(L_{0,i},L_{m,j})$ $\equiv$ $0,\,m\neq n,n\neq 0.$
Thus $f(L_{0,i},L_{m,j})\equiv\lambda[L_{0,i},L_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j})$ for all $m\in\Gamma$ and all $i,j\in\mathbb{Z}$.
(ii) Secondly, we show that
$f(L_{0,i},H_{m,j})\equiv\lambda[L_{0,i},H_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j})$
for all $m\in\Gamma$ and $i,j\in\mathbb{Z}$.
If $m=0$, $H_{0,j}\in C(\mathcal{L}(\Gamma))$, by Lemma 12, $f(L_{0,i},H_{0,j})=0$.
@@
$L_{0,j},H_{m,j}],f(L_{0,j},H_{m,j})]=0,$
then we have $mH_{m,i+j},\sum_{\begin{subarray}{c}n\in\Gamma\\ i,j\in\mathbb{Z}\end{subarray}}(a_{n}L_{n,i+j}+b_{n}H_{n,i+j})]=0$ for all $n\in\Gamma$ and all $i,j\in\mathbb{Z}.$ We conclude $\sum_{\begin{subarray}{c}n\in\Gamma\end{subarray}}-m^{2}a_{n}=0$, thus $a_{n}=0$, then
@@
$nb_{n}H_{n+1,2(i+j)}=-m\lambda H_{m+1,2(i+j)}.$
We have $m=\sum_{\begin{subarray}{c}n\in\Gamma\end{subarray}}n$, which gives $f(L_{0,i},H_{m,j})\equiv\lambda[L_{0,i},H_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j})$ for all $m\in\Gamma$ and all $i,j\in\mathbb{Z}$.
(iii) Thirdly, we show that $f(L_{0,i},x)\equiv\lambda[L_{0,i},x](\mathrm{mod}\mathbb{F}H_{0,i+j})$ for all $x\in\mathcal{L}(\Gamma)$ and $i,j\in\mathbb{Z}$.
@@
Then we have
$f(L_{0,i},x)$ $=$ $\sum_{\begin{subarray}{c}m\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}\left\{a_{m}f(L_{0,i},L_{m,j})+b_{m}f(L_{0,i},H_{m,j})\right\}$
$=$ $\sum_{\begin{subarray}{c}m\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}\left\{a_{m}\lambda[L_{0,i},L_{m,j}]+b_{m}\lambda[L_{0,i},H_{m,j}]\right\}(\mathrm{mod}\mathbb{F}H_{0,i+j})$
@@
It follows that $f(L_{0,i},x)\equiv\lambda[L_{0,i},x](\mathrm{mod}\mathbb{F}H_{0,i+j})$ for all $x\in\mathcal{L}(\Gamma)$ and all $i,j\in\mathbb{Z}$.
(iv)Then, we show that $f(x,y)\equiv\lambda[x,y](\mathrm{mod}\mathbb{F}H_{0,i+j})$ for all $x,y\in\mathcal{L}(\Gamma)$ and all $i,j\in\mathbb{Z}$.
@@
$f(x,y),[L_{0,i},z]]=[[x,y],f(L_{0,i},z)]$
for all $z\in\mathcal{L}(\Gamma),$ we have
$f(x,y)-\lambda[x,y],[L_{0,i},z]]=0$
for all $z\in\mathcal{L}(\Gamma),$ it follows that $f(x,y)-\lambda[x,y]\in C(\mathcal{L}(\Gamma))$, thus $f(x,y)\equiv\lambda[x,y](\mathrm{mod}\mathbb{F}H_{0,i+j})$.
(v)Finally, we show that $f(x,y)\equiv\lambda[x,y]$ for all $x,y\in\mathcal{L}(\Gamma)$.
@@
where $\alpha$ is a bilinear map on $\mathcal{L}(\Gamma)\times\mathcal{L}(\Gamma)\rightarrow\mathbb{F}$. We have
@@
Then
$\alpha([x,y],z)=0$
for all $x,y,z\in\mathcal{L}(\Gamma).$ We conclude $\alpha(x,y)=0$ for all $x,y\in\mathcal{L}(\Gamma)$. Finally, we obtain that $f(x,y)\equiv\lambda[x,y]$ for all $x,y\in\mathcal{L}(\Gamma)$. That is, $f$ is an inner derivation on $\mathcal{L}(\Gamma)$. ∎
@@ 4.2. Applications
If a linear map $\phi$ on Lie algebra $L$, which subjects to $\phi(x),x]=0$ for all $x\in L$, then $\phi$ called a linear commuting map on $L$. If $f(x,y)=[\phi(x),y]=[x,\phi(y)]$, then $f$ is a biderivation on $L$ (see [20]).
@@
A linear map $\phi$ on $\mathcal{L}(\Gamma)$ is commuting if and only if there exist $\lambda\in\mathbb{F}$ and a linear map $\tau:\,\mathcal{L}(\Gamma)\rightarrow C(\mathcal{L}(\Gamma))$ such that $\phi(x)=\lambda x+\tau(x)$ for all $x\in\mathcal{L}(\Gamma)$.
@@
Conversely, let $f(x,y)=[\phi(x),y]$ for all $x,y\in\mathcal{L}(\Gamma)$. Then $f$ is a biderivation on $\mathcal{L}(\Gamma)$. By Theorem 15, we have
$f(x,y)=[\phi(x),y]=\lambda[x,y],\lambda\in\mathbb{F}.$
Since $\phi(x)-\lambda x\in C(\mathcal{L}(\Gamma))$, then we have a linear map $\tau:\mathcal{L}(\Gamma)\rightarrow C(\mathcal{L}(\Gamma))$ such that $\phi(x)=\lambda x+\tau(x)$ for all $x\in\mathcal{L}(\Gamma)$. ∎
Let $(L,[,])$ be a Lie algebra. A commutative postpost-Lie algebra structure on $L$ is a bilinear product $x\cdot y$ on $L$ satisfying the following identifies:
@@
$x\cdot[y,z]=[x\cdot y,z]+[y,x\cdot z],\,\mbox{for all}\,x,y,z\in L.$
We also call $(\mathcal{L}(\Gamma),[,],\cdot)$ is a commutative postpost-Lie algebra (see [17]).
@@
(see [17]) Let $(L,[,],\cdot)$ be a commutative postpost-Lie algebra. If $f:L\times L\rightarrow L$ subjects to $f(x,y)=x\cdot y,\,x,y\in L$, then $f$ is a biderivation on $L$.
@@
Any commutative postpost-Lie algebra structure on $\mathcal{L}(\Gamma)$ is trivial.
@@
Assume $(\mathcal{L}(\Gamma),[,],\cdot)$ is a commutative postpost-Lie algebra. By Theorem 15 and Lemma 17, there exists $\lambda\in\mathbb{F}$ such that
@@
Thus $2\lambda[x,y]=0$, we conclude $\lambda=0$. That is $f(x,y)=x\cdot y=0$. ∎